OpenStudy (anonymous):

MEDAL AND FAN!!!!!!!!! Which of the following expressions is the inverse of the function y equals quantity x plus 5 divided by 7

OpenStudy (shamim):

So ur given equation is y=(x+5)/7

OpenStudy (shamim):

x=?

OpenStudy (shamim):

U know x+5=7y x=7y-5 Right?

OpenStudy (shamim):

But y=f(x) x=f^-1(y)

OpenStudy (shamim):

So f^-1(y)=7y-5 f^-1(x)=?

OpenStudy (shamim):

Just put y=x